# xTrimoPGLM: Unified 100B-Scale Pre-trained Transformer for Deciphering the Language of Protein

### v1

@@ 1.8 Downstream Tasks Datasets & Evaluation Metrics
To systematically evaluate xTrimoPGLM-100B, we have benchmarked 18 downstream protein-related tasks across multiple domains. Table S2 shows a comprehensive overview of our benchmark performance on all evaluated tasks, divided into four main categories: protein structure, protein developability, protein interactions, and protein functions. In this table, ’Struc.’ represents protein structure, ’Dev.’ represents protein developability, ’Inter.’ represents protein interactions, ’Func.’ represents protein functions, and ’Perf.’ denotes performance (%). The $\clubsuit$ denotes the results that we produce or reproduce, while the $\blacklozenge$ represents direct citations from original papers with the same split train/valid/test sets. For any dataset without established benchmarks, we employ the results of our own ESM2-15B with LoRA fine-tuning. The table elucidates these tasks along with the latest SOTA methodologies employed, their respective performances, and the achievements attained by our proposed xTrimoPGLM-100B model. We emphasize that this comparison is primarily from a task-based perspective, where xTrimoPGLM is combined with fine-tuning techniques to achieve the results. The results reveal that xTrimoPGLM-100B significantly outperforms current SOTA approaches in most protein-related tasks, hence catalyzing advancements in this field. Next, we individually delve into these subtasks, elaborating on the corresponding task definitions, dataset processing, evaluation metrics, and other relevant details.
@@
Contact Map. Contact map prediction (Cont. P.) aims to determine whether two residues, i and j, are in contact or not, based on their distance with a certain threshold (<8Å). This task is an important part of the early Alphafold version (?) for structural prediction. The trRosetta dataset (?) is employed and split into 12,041 training samples, 1,505 validation samples, and 1,505 test samples for this task. The evaluation metric used is the Top L/5 accuracy, considering residue pairs with a separation length greater than 6 and a sequence length cutoff of 512.
@@
Fold Classification. Fold class prediction (Fold. P.) is a scientific classification task that assigns protein sequences to one of 1,195 known folds. The dataset employed for this task is based on SCOP 1.75 (?), a release from 2009, and has been widely adopted by DeepSF (?) and Ankh (?). The primary application of this task lies in the identification of novel remote homologs among proteins of interest, such as emerging antibiotic-resistant genes and industrial enzymes (?). The study of protein fold holds great significance in fields like proteomics and structural biology, as it facilitates the analysis of folding patterns, leading to the discovery of remote homologies and advancements in disease research (?).
@@
Secondary Structure. The study of a protein’s secondary structure (Sec. Struc. P.) forms a fundamental cornerstone in understanding its biological function. This secondary structure, comprising helices, strands, and various turns, bestows the protein with a specific three-dimensional configuration, which is critical for the formation of its tertiary structure. In the context of this work, a given protein sequence is classified into three distinct categories, each representing a different structural element: Helix (H), Strand (E), and Coil (C). The datasets applied in this study are originally published by NetSurfP-2.0 (?) and have also been utilized by Ankh (?). The datasets employed for testing in our investigation are specifically assembled from the Critical Assessment of Protein Structure Prediction (CASP) editions 12 and 14, which contain 18 and 21 samples. The result we reported is an average of these two datasets.
@@
Solubility. This task (Sol. P.) involves a binary classification of a heterogenous set of proteins, assessing them as either soluble or insoluble. The solubility metric is a crucial design parameter in ensuring protein efficacy, with particular relevance in the pharmaceutical domain. We’ve adhered to the same dataset division as is employed in the development of DeepSol (?). Within this framework, any protein exhibiting a sequence identity of 30% or greater to any protein within the test subset is eliminated from both the training and evaluation subsets, ensuring robust and unbiased evaluation.
@@
Stability. The task (Stab. P.) is to predict the concentration of protease at which a protein can retain its folded state. Protease, being integral to numerous biological processes, bears significant relevance and a profound comprehension of protein stability during protease interaction can offer immense value, especially in the creation of novel therapeutics. The dataset applied in this task is initially sourced from Rocklin et al (?) and subsequently collected within the Task Assessing Protein Embeddings (TAPE) (?). In this regression-based task, we employ the SpeaRman Correlation Coefficient (SRCC) as the evaluation metric to measure the prediction consistency.
@@
Optimal Temperature. Grasping the catalytic activity of enzymes is pivotal for industrial enzyme design, particularly in predicting the optimal temperature (Opt. Temp.) for a given enzyme’s catalytic effect. The dataset utilized for this task is primarily procured by DeepET (?), a recent advancement in the field that uses deep learning techniques to understand enzyme thermal adaptation. To quantify the relationship between these variables, we use the SRCC.
@@
Optimal PH. Enzyme functions normally under a specific range of pH of the surrounding environment. However, an optimal pH for the reaction will largely boost the catalytic ability. We collected a dataset from EpHod (?), which established a deep learning method to predict the optimal enzyme catalytic pH based on protein sequence only. In the regression-based task, we use the SRCC as the evaluation metric.
@@
Clone CLF. Protein structure determination includes a series of experimental stages to yield stable proteins for X-ray crystallography. Specifically, the proteins are first selected and expressed, then purified for crystal structure determination. Each step corresponds to a "stage tag" to denote whether the protein is stable under a certain stage. We collected a dataset from PredPPCrys (?), which manually annotated thousands of proteins with different experimental procedures. The binary classification task is to predict whether a protein sequence tends to be a cloning failure (Clone CLF.). We use the AUC as the evaluation metric.
@@
Material Production. The task is to predict whether a protein sequence fails at the protein material stage (MF). The dataset is also collected from PredPPCrys (?) and the AUC metric is employed as the measurement.
@@
Metal Ion Binding. Metal ion binding (Metal B.) sites within proteins play a crucial role across a spectrum of processes, spanning from physiological to pathological, toxicological, pharmaceutical, and diagnostic. Consequently, the development of precise and efficient methods to identify and characterize these metal ion binding sites in proteins has become an imperative and intricate task for bioinformatics and structural biology. This task involves a binary classification challenge aimed at predicting the existence of metal-ion binding site(s) on a given protein sequence. We employ data (?) curated from the Protein Data Bank (PDB).
@@
Enzyme Catalytic Efficiency. This task (Enzyme Eff.) is focused on predicting kcat values, which are enzymatic turnover numbers denoting the maximum chemical conversion rate of a reaction, for metabolic enzymes originating from any organism. These predictions are based on substrate structures and protein sequences. The underlying importance of this task lies in its potential to yield high-throughput and accurate kcat predictions applicable to any organism or enzyme. Such capabilities are crucial for advancing our understanding of cellular metabolism and physiology. The data, sourced from a variety of repositories including BRENDA, SABIO-RK, KEGG, UniProt, and MetaCyc, are curated by Li et al (?).
@@
Peptide-HLA/MHC Affinity. The human leukocyte antigen (HLA) gene encodes major histocompatibility complex (MHC) proteins, which can bind to peptide fragments and be presented to the cell surface for subsequent T cell receptors (TCRs) recognition. Accurately predicting the interaction between peptide sequence and HLA molecule will boost the understanding of immune responses, antigen presentation, and designing therapeutic interventions such as peptide-based vaccines or immunotherapies. The classification task aims to predict whether a given paired peptide and HLA sequence can bind or not. The modeling data is from Wu et al (?).
@@
TCR-pMHC Affinity. The interaction between T cell receptors (TCRs) and peptide-major histocompatibility complexes (pMHCs) plays a crucial role in the recognition and activation of T cells in the immune system. TCRs are cell surface receptors found on T cells, and pMHCs are complexes formed by peptides derived from antigens bound to major histocompatibility complexes (MHCs) on the surface of antigen-presenting cells. The classification task is to predict whether a given paired TCR sequence and peptide can bind or not. The evaluated data is major from VDJdb, processed and curated from Pham et al (?).
@@
Antibiotic Resistance. Antibiotic resistance (Antib. Res.) refers to the ability of bacteria and other microorganisms to resist the effects of an antibiotic to which they are once sensitive. In this task (Antib. Res.), an input protein sequence is categorized according to which of 19 antibiotics it is resistant to. Thus, the scope of antibiotic drug development and research could be explored as an understanding in this topic is accumulated. The Dataset used in this task is curated by CARD (?).
@@
Fluorescence. The Fluorescence Prediction (Fluor. P.) task (?) focuses on predicting the fluorescence intensity of green fluorescent protein mutants, a crucial function in biology that allows researchers to infer the presence of proteins within cell lines and living organisms. This regression task utilizes training and evaluation datasets that feature mutants with three or fewer mutations, contrasting the testing dataset, which comprises mutants with four or more mutations. The partitioning of the datasets mirrors the splitting method implemented in the TAPE (?). The quality of these predictions is assessed using the Spearman score as the primary evaluation metric.
@@
Fitness. The task of Fitness Prediction (Fitness P.) is dedicated to anticipating the fitness landscape of the GB1 domain, a process that plays a pivotal role in understanding and enhancing the binding affinity and stability of this domain. As a prevalent protein interaction domain, GB1 finds wide usage in diverse applications such as protein purification, biosensors, and drug delivery (?, ?). This task is configured as a regression problem, where the goal is to predict the fitness of GB1 binding following mutations at four specific positions. The data for this task is sourced from the FLIP database (?). Predictive efficacy is assessed using the Spearman score as the principal evaluation metric.
@@
Localization. The task of Protein Subcellular Localization Prediction (Loc. P.) bears substantial relevance in bioinformatics, owing to its contributions to proteomics research and its potential to augment our comprehension of protein function and disease mechanisms (?). In this task, the input to the model is an amino acid sequence of a protein, which is transformed into an output comprising a probability distribution over 10 unique subcellular localization categories. The dataset applied for this task is derived from Uniprot, meticulously curated by Armenteros et al (?).
@@
We evaluated all benchmarked downstream tasks with xTrimoPGLM-100B and ESM2 models. The performance results are followed as Table S3. Metric values are shown in both probing and LoRA (in parentheses) fine-tuning modes, where the underline denotes the best performance of probing and bold indicates the best performance of LoRA fine-tuning.
@@
1.9 xT-Fold Acceleration Methods
@@
1.10 xT-Fold Training Settings
@@
1.11 The SFT & ReST Pipelines
1.11.1 Dataset
For the comprehensive assessment of the alignment ability of xTrimoPGLM-100B, we curate 5 datasets from the 18 benchmarks described in SM Section 1.8 for supervised fine-tuning and reinforcement self-training.

### v2

@@ 1 Protein Understanding Benchmarks and Evaluations
To systematically evaluate xTrimoPGLM-100B, we have benchmarked 18 downstream protein-related tasks across multiple domains (Supplementary Table 3), which are divided into four main categories: protein structure, protein developability, protein interactions, and protein functions. “Struc.” represents protein structure, “Dev.” represents protein developability, “Inter.” represents protein interactions, “Func.” represents protein functions, and “Perf.” denotes performance (%). The $\clubsuit$ denotes the results that we produce or reproduce, while the $\blacklozenge$ represents direct citations from original papers with the same split train/valid/test sets. For any dataset without established benchmarks, we employ the results of our own ESM2-15B with LoRA fine-tuning. The table elucidates these tasks along with the latest SOTA methodologies employed, their respective performances, and the achievements attained by our proposed xTrimoPGLM-100B model. We emphasize that this comparison is primarily from a task-based perspective, where xTrimoPGLM is combined with fine-tuning techniques to achieve the results. The results reveal that xTrimoPGLM-100B significantly outperforms current SOTA approaches in most protein-related tasks, hence catalyzing advancements in this field. Next, we individually delve into these subtasks, elaborating on the corresponding task definitions, dataset processing, evaluation metrics, and other relevant details.
@@
Contact Map. Contact map prediction (Cont. P.) aims to determine whether two residues, i and j, are in contact or not, based on their distance with a certain threshold (<8Å). This task is an important part of the early Alphafold version  (?) for structural prediction. The trRosetta dataset  (?) is employed and the same split (12,041 training samples, 1,505 validation samples, and 1,505 test samples) as Ankh  (?) is used for this task. The evaluation metric used is the Top L/5 accuracy, considering residue pairs with a separation length greater than 6 and a sequence length cutoff of 512.
@@
Fold Classification. Fold class prediction (Fold. P.) is a scientific classification task that assigns protein sequences to one of 1,195 known folds. The primary application of this task lies in the identification of novel remote homologs among proteins of interest, such as emerging antibiotic-resistant genes and industrial enzymes  (?). The study of protein fold holds great significance in fields like proteomics and structural biology, as it facilitates the analysis of folding patterns, leading to the discovery of remote homologies and advancements in disease research  (?). The dataset employed for this task is based on SCOP 1.75  (?), a release from 2009, and has been widely adopted by DeepSF  (?) and Ankh  (?). We used the same split (12,312 training samples, 736 validation samples, and 3,244 test samples) as Ankh  (?) for evaluation. Total accuracy is employed as the evaluation metric.
@@
Secondary Structure. The study of a protein’s secondary structure (Sec. Struc. P.) forms a fundamental cornerstone in understanding its biological function. This secondary structure, comprising helices, strands, and various turns, bestows the protein with a specific three-dimensional configuration, which is critical for the formation of its tertiary structure. In the context of this work, a given protein sequence is classified into three distinct categories, each representing a different structural element: Helix (H), Strand (E), and Coil (C). The datasets applied in this study are originally published by NetSurfP-2.0  (?) and have also been utilized by Ankh  (?). The training set is the same as Ankh  (?). The datasets employed for testing in our investigation are specifically assembled from the Critical Assessment of Protein Structure Prediction (CASP) editions 12 and 14, which contain 18 and 21 samples. The result we reported is an average of these two datasets. Total accuracy is employed as the evaluation metric.
@@
Solubility. This task (Sol. P.) involves a binary classification of a heterogenous set of proteins, assessing them as either soluble or insoluble. The solubility metric is a crucial design parameter in ensuring protein efficacy, with particular relevance in the pharmaceutical domain. We’ve adhered to the same dataset division as is employed in the development of DeepSol  (?) (62,478 training samples and 6,942 test samples). Within this framework, any protein exhibiting a sequence identity of 30% or greater to any protein within the test subset is eliminated from both the training and evaluation subsets, ensuring robust and unbiased evaluation. Total accuracy is employed as the evaluation metric.
@@
Stability. The task (Stab. P.) is to predict the concentration of protease at which a protein can retain its folded state. Protease, being integral to numerous biological processes, bears significant relevance and a profound comprehension of protein stability during protease interaction can offer immense value, especially in the creation of novel therapeutics. The dataset applied in this task is initially sourced from Rocklin et al  (?) and subsequently collected within the Task Assessing Protein Embeddings (TAPE)  (?). We used the same split (53,614 training samples, 2,512 validation samples and 12,851 test samples) as TAPE  (?) for evaluation. In this regression-based task, we employ the SpeaRman Correlation Coefficient (SRCC) as the evaluation metric to measure the prediction consistency.
@@
Optimal Temperature. Grasping the catalytic activity of enzymes is pivotal for industrial enzyme design, particularly in predicting the optimal temperature (Opt. Temp.) for a given enzyme’s catalytic effect. The dataset utilized for this task is primarily procured by DeepET  (?), a recent advancement in the field that uses deep learning techniques to understand enzyme thermal adaptation. We used the same split (1,706 training samples and 190 test samples) as DeepET  (?) for evaluation. To quantify the relationship between these variables, we use the SRCC.
@@
Optimal PH. Enzyme functions normally under a specific range of pH of the surrounding environment. However, an optimal pH for the reaction will largely boost the catalytic ability. We collected a dataset from EpHod  (?), which established a deep learning method to predict the optimal enzyme catalytic pH based on protein sequence only. We used the same split (7,124 training samples, 760 validation samples and 1,971 test samples) as EpHod  (?) for evaluation. In the regression-based task, we use the SRCC as the evaluation metric.
@@
Clone CLF. Protein structure determination includes a series of experimental stages to yield stable proteins for X-ray crystallography. Specifically, the proteins are first selected and expressed, then purified for crystal structure determination. Each step corresponds to a "stage tag" to denote whether the protein is stable under a certain stage. We collected a dataset from PredPPCrys  (?), which manually annotated thousands of proteins with different experimental procedures. We used the same split (23,375 training samples and 4,791 test samples) as PredPPCrys  (?) for evaluation. The binary classification task is to predict whether a protein sequence tends to be a cloning failure (Clone CLF.). We use the AUC as the evaluation metric.
@@
Material Production. The task is to predict whether a protein sequence fails at the protein material stage (MF). The dataset is also collected from PredPPCrys  (?) (23,339 training samples and 4,791 test samples) and the AUC metric is employed as the measurement.
@@
Metal Ion Binding. Metal ion binding (Metal B.) sites within proteins play a crucial role across a spectrum of processes, spanning from physiological to pathological, toxicological, pharmaceutical, and diagnostic. Consequently, the development of precise and efficient methods to identify and characterize these metal ion binding sites in proteins has become an imperative and intricate task for bioinformatics and structural biology. This task involves a binary classification challenge aimed at predicting the existence of metal-ion binding site(s) on a given protein sequence. We employ data  (?) curated from the Protein Data Bank (PDB). We used the same split (6,000 training samples and 1,332 test samples) for evaluation. Total accuracy is employed as the evaluation metric.
@@
Enzyme Catalytic Efficiency. This task (Enzyme Eff.) is focused on predicting kcat values, which are enzymatic turnover numbers denoting the maximum chemical conversion rate of a reaction, for metabolic enzymes originating from any organism. These predictions are based on substrate structures and protein sequences. The underlying importance of this task lies in its potential to yield high-throughput and accurate kcat predictions applicable to any organism or enzyme. Such capabilities are crucial for advancing our understanding of cellular metabolism and physiology. The data, sourced from a variety of repositories including BRENDA, SABIO-RK, KEGG, UniProt, and MetaCyc, are curated by Li et al   (?). We used the same split (13,470 training samples, 1,684 validation samples and 1,684 test samples) for evaluation. Pearson correlation coefficient (PCC) is employed as the evaluation metric.
@@
Peptide-HLA/MHC Affinity. The human leukocyte antigen (HLA) gene encodes major histocompatibility complex (MHC) proteins, which can bind to peptide fragments and be presented to the cell surface for subsequent T cell receptors (TCRs) recognition. Accurately predicting the interaction between peptide sequence and HLA molecule will boost the understanding of immune responses, antigen presentation, and designing therapeutic interventions such as peptide-based vaccines or immunotherapies. The classification task aims to predict whether a given paired peptide and HLA sequence can bind or not. The modeling data is from Wu et al  (?). The raw dataset contains millions of samples, we used the same split and downsample 1% for training and 5% for validation and testing (57,357 training samples, 7,008 validation samples and 8,406 test samples). The AUC metric is employed as the measurement.
@@
TCR-pMHC Affinity. The interaction between T cell receptors (TCRs) and peptide-major histocompatibility complexes (pMHCs) plays a crucial role in the recognition and activation of T cells in the immune system. TCRs are cell surface receptors found on T cells, and pMHCs are complexes formed by peptides derived from antigens bound to major histocompatibility complexes (MHCs) on the surface of antigen-presenting cells. The classification task is to predict whether a given paired TCR sequence and peptide can bind or not. The evaluated data is major from VDJdb, processed and curated from Pham et al  (?). We used the vdjdb_no10x split (19,526 training samples and 4,485 test samples) for evaluation. The AUC metric is employed as the measurement.
@@
Antibiotic Resistance. Antibiotic resistance (Antib. Res.) refers to the ability of bacteria and other microorganisms to resist the effects of an antibiotic to which they are once sensitive. In this task (Antib. Res.), an input protein sequence is categorized according to which of 19 antibiotics it is resistant to. Thus, the scope of antibiotic drug development and research could be explored as an understanding in this topic is accumulated. The Dataset used in this task is curated by CARD  (?). We used the same split (2,072 training samples and 1,344 test samples) for evaluation. Total accuracy is employed as the evaluation metric.
@@
Fluorescence. The Fluorescence Prediction (Fluor. P.) task  (?) focuses on predicting the fluorescence intensity of green fluorescent protein mutants, a crucial function in biology that allows researchers to infer the presence of proteins within cell lines and living organisms. This regression task utilizes training and evaluation datasets that feature mutants with three or fewer mutations, contrasting the testing dataset, which comprises mutants with four or more mutations. The partitioning of the datasets mirrors the splitting method implemented in the TAPE  (?)(21,446 training samples, 5,362 validation samples and 27,217 test samples). The quality of these predictions is assessed using the Spearman score as the primary evaluation metric.
@@
Fitness (GB1). The task of Fitness Prediction (Fitness P.) is dedicated to anticipating the fitness landscape of the GB1 domain, a process that plays a pivotal role in understanding and enhancing the binding affinity and stability of this domain. As a prevalent protein interaction domain, GB1 finds wide usage in diverse applications such as protein purification, biosensors, and drug delivery  (?). This task is configured as a regression problem, where the goal is to predict the fitness of GB1 binding following mutations at four specific positions. The data for this task is sourced from the FLIP database  (?) and the sampled split is used (6,289 training samples, 699 validation samples and 1,745 test samples). Predictive efficacy is assessed using the Spearman score as the principal evaluation metric.
Zero-shot Fitness Prediction. We utilize the DMS_ProteinGym_substitutions dataset from the ProteinGym benchmark  (?) to evaluate zero-shot protein fitness predictions. Due to the ESM models’ length constraints, we filtered out sequences exceeding 1024 residues, resulting in 201 assays with over 8 million mutations. For protein understanding models like the ESM family, we use the metric from ESM-1v  (?) to score mutations based on the log odds ratio at the mutated positions. In sequences with multiple mutations, we assume an additive model. Mutation scores are calculated using the wild type and mutated type, with each mutated position masked in the sequence input. For generative models like PROGEN2  (?), we use the log-likelihood of the entire mutated sequence as the predicted mutation score. Performance is assessed by determining the Spearman correlation coefficient between predicted and ground truth scores. Results are averaged within each assay and then across all assays. For comprehensive comparison, we also include previous supervised fitness predictions on the Fitness GB1 dataset, as shown in Supplementary Table 7. We observe that the scaling effect does not seem to hold in zero-shot fitness prediction, which aligns with observations from other works like PROGEN2. We hypothesize that although large pre-trained models may have a stronger capacity to differentiate between original and mutated sequences, this ability might not directly translate to accurate mutation scores without supervision. Thus, our xT100B model, after SFT, still holds better performance compared with other smaller models on the supervised fitness prediction scenarios.
@@
Localization. The task of Protein Subcellular Localization Prediction (Loc. P.) bears substantial relevance in bioinformatics, owing to its contributions to proteomics research and its potential to augment our comprehension of protein function and disease mechanisms  (?). In this task, the input to the model is an amino acid sequence of a protein, which is transformed into an output comprising a probability distribution over 10 unique subcellular localization categories. The dataset applied for this task is derived from Uniprot, meticulously curated by Armenteros et al  (?) (6,622 training samples and 1,842 test samples). Total accuracy is employed as the evaluation metric.
@@
We evaluated all benchmarked downstream tasks with xTrimoPGLM-100B and ESM2 models (Supplementary Table 4). Metric values are shown in both probing and LoRA (in parentheses) fine-tuning modes, where the underline denotes the best performance of probing and bold indicates the best performance of LoRA fine-tuning.
@@
2 xT-Fold Acceleration Methods
@@
3 xT-Fold Training Settings
@@
Below, we provide additional methodological insights into the architecture and training procedures of xT-Fold.
Model Architecture: The key modification involves employing MSARowAttentionWithPairBias exclusively for the input query sequence, rather than attending to each sequence individually in MSAs, and we further exclude the MSAColumnAttention block. We employ a 48-layer version of this block with up to three recycling stages during training and utilize $\text{recycle}_{\text{num}}=3$ during inference. This framework is a refined adaptation based on AF2’s Evoformer framework. Despite minor adjustments, we retain the Evoformer terminology, but it more closely resembles a simplified version within the current AF3 framework, referred to as Pairformer  (?). This adjustment significantly enhances training and inference speeds.
Initialization: During initialization, we treat the Evoformer as a residual block by initializing its output parameters to zero. This initialization establishes a quasi-identity mapping from PLM to the structure module, leveraging Evoformer’s residual modules. Throughout the training, Evoformer modules are dynamically selected based on gradients. We froze the entire PLM during training without fine-tuning. Hyperparameters: Hyperparameter settings closely follow those of AF2 and ESMFold, including: Exponential Moving Average (EMA): Decay 0.999; Adam Optimizer Parameters: Base learning rate: 1e-3, $\beta_{1}$ = 0.9, $\beta_{2}$ = 0.999, eps = 1e-8.
Data Distillation: For data distillation in training, we initially filtered approximately 60 million samples from AFDB based on AF2’s pLDDT 0.9 criteria, resulting in approximately 4 million samples after clustering with a sequence max identity of 0.5. Subsequently, we filtered out parts similar to the test set by 90%, and integrated 38K samples from PDB in a 75:25 (#Distill:#PDB) ratio blend. Throughout this process, we employed cluster sampling based on sequence length and cluster size, with a sampling probability calculated as:
[…]
Training Procedure: The training process consists of two stages: Stage 1: Trained on 10 million samples with Crop size: 256 and Batch size: 128; Stage 2: Increased crop size to 384. Incorporated pLDDT and a minimally weighted violent loss (coefficient set to 0.01) Trained on an additional 1 million samples. Used cosine decay with the learning rate decreasing from 5e-5 to 1e-5.
4 AlphaFold2 with FlashAttention
Since we hope to use FlashAttention to speed up the running of AlphaFold2, we use OpenFold’s AlphaFold2 implementation code to load DeepMind’s AlphaFold2 weights (by –jax_param_path option) for inference. In our previous tests, we have confirmed that structure prediction by using OpenFold code + Jax weight and directly using DeepMind’s official version for inference have exactly the same results. FlashAttention significantly reduces the time and memory consumption of original Self-Attention by reducing the number of reads and writes between high bandwidth memory (HBM) and GPU on-chip SRAM, without changing the results of numerical computation.
For each query sequence, we use DeepMind’s official MSA retrieval pipeline to obtain the MSA sequences and PDB templates and limited the maximum PDB template release date to 2020-05-01. We used five checkpoints of AF2 to run structure prediction. For each checkpoint, we set different random number seeds and run five times, obtaining a total of 25 predicted models (without AMBER relaxation). We calculated the TM-score and GDT_TS score of all models and found that there was little difference between all models (CAMEO TM-score 0.87 ~0.88, CASP15 TM-score 0.74 ~0.77). It should be noted that AlphaFold2’s model_4 and model_5 have PDB templates as input, and our xT-Fold doesn’t. For a more fair comparison, we use the first run result of model_3 (model_3_0) as the representative result of AlphaFold2.
5 The SFT & ReST Pipelines
5.1 Dataset
For the comprehensive assessment of the alignment ability of xTrimoPGLM-100B, we curate 5 datasets from the 18 benchmarks for supervised fine-tuning and reinforcement self-training (Supplementary Section 1).